The condensed income statements for years 20x1 and 20x2 for the Ruhf Inn are shown below:

20x1 20x2

Room revenue $ 2,500,000 $3,000,000

Room labor costs 400,000 500,000

Room other expenses 300,000 350,000

Room income 1,800,000 2,150,000

Undistributed operating expenses 500,000   600,000

Interest expense 200,000 250,000

Rent expense 100,000 120,000

Other fixed charges 400,000 400,000

Income before taxes 600,000 780,000

Income taxes 200,000 260,000

Net Income 400,000 520,000

Required:

1. Determine the following for each year:

a. Number of times interest earned ratio

b. Fixed charge coverage ratio

c. profit margin

d. Room labor cost percentage

e. Average tax rate

2. Prepare a common size income statement for 20x1 and 20x2 and comment on any significant differences in the percentages
Answer #1

1.

No of time interest earned ratio EBIT/Interest (600,000+200,000)/200,000 = 4 times (780000+250000)/400000 = 4.12 times  
Fixed charge coverage ratio EBIT+fixed charge/Interest+fixed charge

(800000+400000+100000)/(200,000+400,000+100,000) =

1.86 times

(1030000+400000+120000)/(250000+400000+120000) =

2.01 times

Profit margin (assuming net profit) Net Income/Room Revenue*100 400000/2500000*100 = 16% 520000/3000000*100 =17.33%
Room Labor cost % (assuming % of sales) Labor Cost/room revenue*100 400000/2500000*100 = 16% 500000/3000000*100 = 16.67%

2.  

Common size statement
20x1 20x2
Room revenue    2,500,000.00 100.00%    3,000,000.00 100.00%
Room labor cost        400,000.00 16.00%        500,000.00 16.67%
Room other expenses        300,000.00 12.00%        350,000.00 11.67%
Gross profit    1,800,000.00 72.00%    2,150,000.00 71.67%
Undistributed Operating expense        500,000.00 20.00%        600,000.00 20.00%
Rent expense        100,000.00 4.00%        120,000.00 4.00%
other fixed charges        400,000.00 16.00%        400,000.00 13.33%
Operating income (EBIT)        800,000.00 32.00%    1,030,000.00 34.33%
Interest        200,000.00 8.00%        250,000.00 8.33%
EBT        600,000.00 24.00%        780,000.00 26.00%
Taxes thereon        200,000.00 8.00%        260,000.00 8.67%
Net income        400,000.00 16.00%        520,000.00 17.33%